OpenStudy (anonymous):

If the integral from 0 to 3 of f(x)dx=6 and the integral from 3 to 5 of f(x)=4, then the integral of 5 to 0 of (3+2f(x))dx= ... The answer is 35. Please help how to get to this answer.

OpenStudy (anonymous):

\[ \int\limits_5^03+2f(x)\;dx = 3\int\limits_5^0\;dx+2\int\limits_5^0f(x)\;dx \]

OpenStudy (anonymous):

You should be able to solve\[ \int\limits_5^0dx \]Relatively simply.

OpenStudy (anonymous):

Notice that: \[ \int \limits_5^0 f(x)\;dx = -\int \limits_0^5 f(x)\;dx = -\left(\int \limits_0^3 f(x)\;dx+\int \limits_3^5 f(x)\;dx\right) \]

OpenStudy (anonymous):

I got it. :) thanks for your help!